OpenStudy (anonymous):

prove if 3 | (a^2+b^2) then 3 | a and 3 | b

Parth (parthkohli):

\[3n = a^2 + b^2 \iff 3n - a^2 = b^2 \iff 3\left(n-\dfrac{a^2}{3}\right) = b^2\]This means that \(a^2\) must be divisible by \(3\) (get why?) and if so, then \(b^2\) is also a multiple of \(3\). If \(3|a^2\) then \(3|a\) and \(3|b^2\) then \(3|b\). Q.E.D.

OpenStudy (anonymous):

what does Q.E.D. mean?

Parth (parthkohli):

Proved.

OpenStudy (anonymous):

thank you

Parth (parthkohli):

You're welcome, I hope you get what my proof says :-)

OpenStudy (anonymous):

I appreciate your work and if I wasn't able to read mathematical proofs I would not have made it this far.

Parth (parthkohli):

Yay!